OpenStudy (anonymous):

Solve the system algebraically. 2x + y - 10 = 0 x - y - 5 = 0 What is the value of y?

OpenStudy (amistre64):

let x=2+n let y=-3+n

OpenStudy (amistre64):

plug in and solve for n to determine x and y

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

whereever you see an x or y, plug in the values for them that have been defined. then you only have to determine the solution of 1 variable instead of 2 ....

OpenStudy (amistre64):

2x + y - 10 = 0 x - y - 5 = 0 2(2+n) + (-3+n) - 10 = 0 (2+n) - (-3+n) - 5 = 0 solve for n